OpenStudy (anonymous):

PLS HELP: Give a value of A so that the function f(x)= {x-x^2, x<2 {x^2 - Ax, x>=2 is continuous. Usually I check to see if the f(x) exists and then go from there, but the A is in that equation!

OpenStudy (anonymous):

Continuous means lim\[\lim_{x \rightarrow a}f(x)=f(a)\] So we just plug in 2. 2-2^2=2^2-2A -2=4-2A 2A=6 A=3

OpenStudy (anonymous):

wow, you're good!

OpenStudy (anonymous):

that really helped me understand it- thanks a ton!

OpenStudy (anonymous):

so you made the left hand side equal to the right hand side as both must agree, as well as f(x) for the it to be continuous?

OpenStudy (anonymous):

Right, the limits from both sides must be equal, so I just plugged 2 in (remember that if you plug a number into a function, it'll give you the limit, provided that you get something normal...i.e. not divided by 0, etc.).

OpenStudy (anonymous):

ok, gotcha!

myininaya (myininaya):

gj goblue